### v1

# CAF-I: A Collaborative Multi-Agent Framework for Enhanced Irony Detection with Large Language Models

Abstract
Large language model (LLM) have become mainstream methods in the field of sarcasm detection. However, existing LLM methods face challenges in irony detection, including: 1. single-perspective limitations, 2. insufficient comprehensive understanding, and 3. lack of interpretability. This paper introduces the Collaborative Agent Framework for Irony (CAF-I), an LLM-driven multi-agent system designed to overcome these issues. CAF-I employs specialized agents for Context, Semantics, and Rhetoric, which perform multidimensional analysis and engage in interactive collaborative optimization. A Decision Agent then consolidates these perspectives, with a Refinement Evaluator Agent providing conditional feedback for optimization. Experiments on benchmark datasets establish CAF-I’s state-of-the-art zero-shot performance. Achieving SOTA on the vast majority of metrics, CAF-I reaches an average Macro-F1 of 76.31%, a 4.98% absolute improvement over the strongest prior baseline. This success is attained by its effective simulation of human-like multi-perspective analysis, enhancing detection accuracy and interpretability.
Multi-Agent System, Irony Detection, Large Language Model
1 Introduction
The rapid development of social media has fostered increasingly diverse and semantically complex forms of textual expression (Kader et al., 2022) highlights the challenge of irony detection. Irony, a rhetorical device contrasting stated and intended meaning (Booth, 1974), is crucial for NLP systems as its presence alters textual interpretation, impacting applications like sentiment analysis and content moderation (Reyes et al., 2013). However, detection is complicated by context dependency, obscure authorial intent, and varied rhetorical strategies (Wallace et al., 2015).
Initial efforts used traditional machine learning with handcrafted features (Reyes et al., 2013), often failing to capture irony’s subtleties. Deep learning methods improved semantic representation but struggled with the implicit and contextual nature of ironic utterances (Ghosh & Veale, 2016; Baziotis et al., 2018). Consequently, the emergence of LLM has presented novel avenues, exhibiting substantial potential in irony detection, typically via fine-tuning or advanced prompt engineering techniques (Yao et al., 2024).
Despite this significant promise and their advanced capabilities, LLM encounters a distinct set of challenges when specifically applied to the nuanced task of irony detection, as illustrated in Figure 1. Key among these challenges are:
  1. 1.
    Single-Perspective Limitation: As single-model predictors, LLM have limited capacity for multidimensional collaborative reasoning, struggling to collaboratively synthesize diverse analytical insights required for deconstructing complex irony.
  2. 2.
    Insufficient Comprehensive Understanding: LLM struggle to holistically integrate diverse informational cues, such as context, semantics, and rhetoric, to achieve human-like deep comprehension of ironic intent.
  3. 3.
    Lack of Interpretability: Opaque reasoning process in irony judgments hinders the understanding, trust, and debugging of decision-making mechanisms.
Figure 1: Examples illustrating three primary challenges for LLM in sarcasm detection: single perspective limitation, insufficient comprehensive understanding, and lack of interpretability.
Motivated by Chain-of-Thought (CoT) prompting (Wei et al., 2022) and multi-agent cooperation principles (Liu et al., 2023), this paper introduces CAF-I, a novel LLM-driven multi-agent framework to address existing limitations in irony detection by simulating human-like multidimensional analytical processes. CAF-I initially performs multi-perspective parallel analysis from three core angles: contextual consistency, deep semantic logic, and rhetorical function. Subsequently, an interactive collaborative optimization mechanism facilitates information sharing and debate among these perspectives. Finally, an integrated decision mechanism allows a central decision-making agent to consolidate these viewpoints and apply conditional optimization triggered by evaluative feedback, forming the paper’s primary contributions:
  • Introduces CAF-I, a novel LLM-driven multi-agent framework significantly improving irony detection through specialized agent collaboration and phased optimization.
  • Achieves state-of-the-art zero-shot performance across multiple benchmark datasets.
  • Its unique architecture offers clear decision pathways and error correction capabilities, enhancing interpretability.
  • Extensive experiments validate its component necessity, overall robustness, and innovation, establishing CAF-I as a pioneering application of specialized LLM agents in structured collaboration for irony detection.
2 RELATED WORK
2.1 Irony Detection
Irony detection in NLP remains challenging due to its inherent complexity, context dependence, and contradictory nature. Early machine learning relied on hand-crafted features, including lexical cues (Davidov et al., 2010) and sentiment lexicons (Reyes et al., 2013), but inadequately captured irony’s subtleties. Deep learning techniques offered significant improvements: Word embeddings such as Word2Vec (Mikolov et al., 2013) and GloVe (Pennington et al., 2014) enabled richer semantic representations. Subsequently, Convolutional Neural Networks (Poria et al., 2016), recurrent neural networks, and particularly LSTM architectures (Zhang et al., 2016) learned hierarchical and sequential features, while Graph Neural Networks (Liang et al., 2022) captured structural dependencies to improve performance. Nonetheless, these deep learning approaches still faced difficulties with irony’s implicit meanings and complex rhetoric, motivating the exploration of more advanced models.
Recent advances in large language models have introduced prompt-based and zero-shot learning methods, allowing models like GPT-4o to interpret irony with minimal reliance on labeled data (Yao et al., 2024). By leveraging task-specific prompts, these models have demonstrated enhanced context-aware irony classification. However, existing LLM-based approaches still operate as single-model predictors, lacking the capacity for multi-perspective reasoning.
2.2 CoT Prompt Engineering
To address the limitations of direct predictions in complex reasoning tasks, CoT prompting emerged as a structured technique that enhances language models by guiding them through intermediate reasoning steps (Wei et al., 2022). Despite its effectiveness, CoT initially depended on manually created prompts, which constrained its adaptability. Subsequent innovations, such as Auto-CoT, addressed this by automating the creation of reasoning chains (Zhang et al., 2022). Further extensions, such as Tree-of-Thought (ToT), expanded reasoning by enabling the exploration of multiple pathways (Yao et al., 2024), while Graph-of-Thought (GoT) introduced a structured approach where reasoning steps are represented as interconnected nodes in a graph (Besta et al., 2024). These advancements directly influenced our multi-agent framework, where specialized agents leverage structured reasoning to collaboratively enhance irony detection.
2.3 Multi-agent Cooperation
Multi-agent frameworks effectively leverage collaborative interactions among specialized agents for complex tasks, with studies exploring forms like deliberation, structured debates, and dialogues. Debating frameworks, for instance, improve factual accuracy and solution diversity in complex reasoning (Du et al., 2023). Others, such as CAMEL (Li et al., 2023), use role-play to simulate nuanced human behaviors; MathChat (Wu et al., 2023b) employs structured dialogues for intricate tasks; and AutoGen (Wu et al., 2023a) shows multi-agent versatility through customizable structures. Inspired by these methodologies, our framework uses specialized LLM-based agents for contextual, semantic, and rhetorical analysis via CoT reasoning. A final decision agent integrates their outputs, enhancing irony detection’s robustness, accuracy, and interpretability.
3 Methodology
This section details our methodology for irony detection, centered on a novel multi-agent collaborative framework driven by LLM, termed the CAF-I. This framework comprises a Context Agent, a Semantic Agent, a Rhetoric Agent, a Decision Agent, and a Refinement Evaluator Agent.
CAF-I aims to address single-model limitations in irony detection via multi-agent collaboration. As irony comprehension demands nuanced understanding of context, semantics, and rhetoric, CAF-I employs dedicated agents for these dimensions. A Decision Agent aggregates these analyses, refined by the conditional feedback of the refinement evaluator agent, to achieve robust and accurate detection.
3.1 Problem Definition
[…]
Figure 2: The Overall architecture of CAF-I. The framework includes three specialized agents, Context, Semantic, and Rhetoric, which conduct multi-round analysis and optimization of the input text. The Decision Agent integrates the preliminary classification, and the Evaluator Agent determines if a round of feedback optimization is needed.
3.2 Overall Workflow
The CAF-I inference workflow (Figure 2) unfolds in key stages: Context (CA), Semantic (SA), and Rhetoric (RA) agents first independently provide initial assessments and reasoning. These are shared for collaborative reanalysis where agents refine judgments using peer insights. A Decision Agent (DA) then aggregates these refined judgments for an initial irony classification. Finally, a Refinement Evaluator Agent (REAgent) reviews this, potentially triggering a single conditional refinement loop with targeted feedback to improve the final classification within one iteration. The detailed design is presented in 3.3.
3.3 Agent Design
3.3.1 Context Agent
(1) Role
The Context Agent identifies irony via semantic analysis of input $x_{i}$ and strategic external knowledge use, detecting inconsistencies between literal and contextual meanings.
(2) Workflow
Upon receiving an input text $x_{i}$, the Context Agent extracts internal contextual features using an LLM. This involves identifying key named entities $E=\{e_{1},...,e_{k}\}$, their semantic relationships $R=\{r_{1},...,r_{l}\}$, and the overarching theme $T$. These components collectively define the internal context, denoted as $\mathcal{C}_{int}=(E,R,T)$. The overall extraction process can be formulated as:
$\mathcal{C}_{int}=\text{ExtractInternalContext}{\text{LLM}}(x_{i})$ (1)
Subsequently, the CA, guided by an LLM, assesses the need for external knowledge based on the ambiguity of $x_{i}$ and the novelty of $\mathcal{C}_{int}$. If deemed beneficial, a search query $q$ is formulated using $q=g_{\text{LLM}}(\text{keywords}(x_{i}))$. Retrieval via an external search API yields a set of documents $D=\{d_{1},...,d_{p}\}$. These are summarized into an external context representation $S$.
[…]
$(y_{1},R_{1})=\text{IronyClassifier}_{\text{LLM}}(x_{i},E,R,T,S)$ (2)
3.3.2 Semantic Agent
(1) Role
The Semantic Agent employs a CoT process for detailed semantic analysis to identify irony. It focuses on the interplay of emotional tone, implied meaning, and commonsense expectations to detect nuanced incongruities.
(2) Workflow
The workflow initiates with an LLM performing semantic parsing of the input $x_{i}$ to discern its literal meaning $M_{lit}(x_{i})$ from pragmatically implied intent $I_{imp}(x_{i})$:
$(M_{lit}(x_{i}),I_{imp}(x_{i}))=\text{SemanticParse}_{\text{LLM}}(x_{i})$ (3)
Subsequently, as part of its CoT reasoning, the agent identifies the expressed emotion $E_{exp}(x_{i})$:
$E_{exp}(x_{i})=\text{EmotionClassifier}_{\text{LLM}}(x_{i})$ (4)
It then evaluates the consistency between $E_{exp}(x_{i})$ and the contextually expected emotion $E_{ctx}(x_{i})$. Concurrently, the SA assesses the text’s consistency with commonsense knowledge $\mathcal{K}_{cs}$:
$\text{Consistency}(x_{i},\mathcal{K}_{cs})$ $=\text{Verifier}_{LLM}(x_{i},\mathcal{K}_{cs})$ (5)
$\in\{\text{Consistent},\text{Inconsistent}\}.$
The SA’s final irony classification $y_{2}$ and its supporting explanation $R_{2}$ are generated by an LLM that holistically reasons over the outcomes of semantic parsing, the evaluation of emotional consistency, and commonsense reasoning through a CoT process:
[…]
$\qquad E_{exp}(x_{i}),E_{ctx}(x_{i}),\text{Consistency}(x_{i},\mathcal{K}_{cs}))$
3.3.3 Rhetoric Agent
(1) Role
The Rhetoric Agent identifies irony by analyzing rhetorical devices within the input text $x_{i}$. It evaluates their interplay with communicative intent, using CoT prompting, to assess irony based on linguistic form.
(2) Workflow
Employing an LLM with CoT, the RA first identifies rhetorical devices $D_{r}$ present in the text:
$D_{r}=\text{RhetoricalDeviceDetector}_{\text{LLM}}(x_{i})=\{d_{r1},...,d_{rm}\}$ (7)
For each detected device $d_{rj}\in D_{r}$, the LLM explains its rhetorical function $E_{rj}$:
$E_{rj}=\text{ExplainRhetoricalFunction}_{\text{LLM}}(d_{rj})$ (8)
Subsequently, the RA then assesses if the overall rhetorical structure, guided by $D_{r}$ and their explanations $\{E_{rj}\}_{j=1}^{m}$, indicates irony, potentially evidenced by contradictions, reversals, or exaggerations. An LLM-based classifier generates the final judgment $y_{3}$ and explanation $R_{3}$ from these inputs:
$(y_{3},R_{3})=\text{RhetoricalIronyClassifier}_{\text{LLM}}(D_{r},\{E_{rj}\}_{j=1}^{m})$ (9)
3.3.4 Decision Agent
(1) Role
The final Decision Agent produces a robust and interpretable irony classification $y$ by consolidating the analysis agents’ second-round outputs, comprising their judgments and reasoning traces.
(2) Workflow
The Decision Agent initiates its process by collecting the second-round outputs from the three primary analysis agents: their binary judgments $y_{1}^{(2)},y_{2}^{(2)},y_{3}^{(2)}$ and their corresponding reasoning traces $R_{1}^{(2)},R_{2}^{(2)},R_{3}^{(2)}$.
The DA employs a hierarchical strategy. If agents achieve consensus, their unanimous judgment, denoted as $y^{*}$, is adopted:
$y=y^{*},\quad\text{if }y_{1}^{(2)}=y_{2}^{(2)}=y_{3}^{(2)}$ (10)
If a two-agent majority exists, the decision is made by majority vote:
$y$ $=\text{majority}(y_{1}^{(2)},y_{2}^{(2)},y_{3}^{(2)}),\quad\text{if }(y_{1}^{(2)}=y_{2}^{(2)}\neq y_{3}^{(2)})$ (11)
$\qquad\lor(y_{1}^{(2)}=y_{3}^{(2)}\neq y_{2}^{(2)})\lor(y_{2}^{(2)}=y_{3}^{(2)}\neq y_{1}^{(2)})$
In cases of complete disagreement, an LLM analyzes the agents’ reasoning traces ($R_{i}^{(2)}$) for clarity, coherence, and relevance, adopting the judgment supported by the most compelling argument.
[…]
3.3.5 Refinement Evaluator Agent
(1) Role.
The Refinement Evaluator Agent operates subsequently, once per inference, to assess the preliminary decision $s$ and explanation $\mathcal{E}$ from the Decision Agent. It determines if a single refinement iteration is necessary ($R_{needed}$) and generates hypothetical feedback as a triplet $(f_{CA},f_{SA},f_{RA})$ for upstream agents to improve reliability.
(2) Workflow.
[…]
$R_{needed}=\begin{cases}\text{"true"}&\text{if }C_{RE}\in\{\text{Low}\}\lor F_{contra}=\text{Yes}\\ \text{"false"}&\text{otherwise}\end{cases}$ (12)
[…]
4 EXPERIMENT
In this section, we try to answer the following research questions:
  • RQ1: How does CAF-I perform on benchmark datasets against diverse existing irony detection methods?
  • RQ2: What are the performance contributions of CAF-I’s core components?
  • RQ3: How robust is CAF-I’s architecture with different LLM backbones, and how does it compare to standard prompting of these LLM?
  • RQ4: How interpretable is CAF-I, and how valuable are its agents’ intermediate reasoning steps for detection performance?
  • RQ5: How does CAF-I’s inference efficiency compare to other competitive CoT-based LLM approaches?
4.1 Experiment Setup
To evaluate the efficacy and generalizability of CAF-I, we utilized four established sarcasm detection benchmarks: IAC-V1, IAC-V2, MUSTARD, and SemEval-2018 Task 3. Model performance was assessed using Accuracy and the Macro-F1 score, with Macro-F1 serving as the primary evaluation metric. The GPT-40 model was employed as the LLM backbone for all agents within the CAF-I framework, and a temperature setting of 0 was used for all API interactions to ensure reproducibility. We compared CAF-I against a comprehensive suite of baselines, including LLM-based approaches, fine-tuned Pre-trained Language Models, and deep learning method. For complete details regarding the experimental setup, including specific dataset characteristics, a detailed description of evaluation metrics, implementation specifics, and a comprehensive list of comparison baselines, please refer to Appendix A.
4.2 Overall Performance Comparison
Table 1: Overall performance comparison across four benchmark datasets. All LLM strategies are zero-shot. Acc. denotes Accuracy and Ma-F1 signifies Macro-F1. Best results are presented in bold, second-best are underlined. Scores are reported as %.
Method IAC-V1 IAC-V2 MuSTARD SemEval-2018 Avg.
Acc. Ma-F1 Acc. Ma-F1 Acc. Ma-F1 Acc. Ma-F1 Acc. Ma-F1
MIARN (Tay et al., 2018) 63.21 63.18 72.75 72.75 64.60 63.90 68.50 67.80 67.26 66.91
SAWS (Pan et al., 2020) 66.13 65.60 76.20 76.20 69.71 70.95 69.90 68.90 70.48 70.41
DC-Net (Liu et al., 2021) 66.50 66.40 78.00 77.90 71.28 71.43 70.80 69.60 71.64 71.33
BERT (Devlin et al., 2019) 65.30 65.20 76.40 76.20 64.30 64.30 69.90 68.40 68.97 68.52
RoBERTa (Liu et al., 2019) 70.10 69.90 76.60 76.70 66.10 66.00 70.20 69.10 70.75 70.42
GPT-4o (Yao et al., 2024) 70.63 70.05 73.03 71.99 67.24 65.79 64.03 63.17 68.73 67.75
GPT-4o+CoC (Yao et al., 2024) 72.19 71.52 73.36 72.31 69.42 68.48 70.79 70.60 71.44 70.73
GPT-4o+Goc(Yao et al., 2024) 65.00 62.91 64.97 61.30 70.69 69.91 74.03 74.02 68.67 67.04
GPT-4o+Boc(Yao et al., 2024) 68.75 67.36 71.35 69.39 69.42 68.45 62.12 61.85 67.91 66.76
CAF-I (Ours) 73.75 73.71 77.87 76.82 75.21 74.73 80.73 79.99 76.89 76.31
The overall performance comparison, with results detailed in table 1, reveals several key insights.
Our Framework Achieves New SOTA. Our proposed CAF-I framework demonstrates clear superiority, establishing a new state-of-the-art. It achieves the highest average Accuracy of 76.89 percent and an average Macro-F1 of 76.31 percent across all benchmarks, consistently outperforming other methods on most datasets and metrics. This underscores the efficacy of CAF-I’s structured multi-agent collaborative reasoning for irony detection.
CoT in CAF-I Outperforms Simpler Prompting. Within LLM-based approaches, our CAF-I framework clearly demonstrates the significant advantage of its CoT reasoning over simpler zero-shot prompting. Specifically, CAF-I’s multi-agent CoT process achieves an average Macro-F1 of 76.31 percent, substantially outperforming the GPT-4o zero-shot baseline’s average Macro-F1 of 67.75 percent. This highlights the substantial benefits derived from employing a sophisticated CoT methodology like that inherent in CAF-I.
LLM-based Methods Outperform Non-LLM Approaches. Finally, LLM-based methods, on the whole, exhibit stronger performance compared to non-LLM approaches. Advanced LLM systems like CAF-I, with an average Macro-F1 of 76.31 percent, substantially outperform leading non-LLM methods; for example, the fine-tuned RoBERTa achieved an average Macro-F1 of 70.42 percent, and the traditional deep learning model DC-Net reached 71.33 percent. This indicates the enhanced capability of sophisticated LLM architectures for complex irony understanding over earlier paradigms.
4.3 Ablation Study
To assess the contribution of each core component within CAF-I, we conducted ablation studies, with detailed results presented in Figure 3. These studies involved systematically removing specialized analysis agents CA, SA, or RA from all analysis stages, or deactivating the REAgent’s conditional refinement mechanism, thereby making the initial aggregated decision final. Experiments were performed on the IAC-V1, MuSTARD, and SemEval-2018 datasets.
The results affirm the integral role of every component. Removing any single analysis agent consistently led to significant performance degradation across all datasets, underscoring their critical contributions. For instance, ablating the RA typically incurred a substantial average drop in Macro-F1. Similarly, deactivating the REAgent’s refinement mechanism also noticeably reduced performance, particularly on more challenging datasets like SemEval-2018, confirming the value of this step for enhancing decision robustness. These findings validate the necessity of each specialized agent and the refinement process within CAF-I.
Figure 3: The ablation experimnet result of CA, SA, RA and REAgent.
4.4 Robust Experiment
To assess CAF-I’s architectural robustness, we evaluated its performance using alternative LLM backbones, specifically Qwen 2-7B and Llama 3-8B, on the four benchmark datasets while maintaining the core framework and prompts. These results were compared against each LLM’s baseline IO prompt performance, sourced from SarcasmCue (Yao et al., 2024), to isolate architectural contributions. The evaluation confirmed CAF-I’s architectural value. Across different LLM backbones, CAF-I consistently demonstrated substantial average Macro-F1 improvements over their respective IO prompting baselines. For instance, CAF-I with Qwen 2-7B achieved a notable 81.25% Macro-F1 on SemEval-2018, surpassing even its primary GPT-4o backbone implementation 79.99% Ma-F1 on this dataset. This general trend of outperformance strongly confirms CAF-I’s robustness and broad applicability. Detailed comparative results are presented in Appendix C.
4.5 Explainability Experiment
4.5.1 Case Study
To illustrate the interpretability and error correction capabilities of CAF-I, a case study was conducted. In this instance (detailes are provided in Figure 4), CAF-I initially an incorrect ”Sarcasm” judgment arose because the Semantic and Rhetoric Agents over-interpreted certain phrases, despite the Context Agent finding no significant inconsistency. However, the REAgent identified this discrepancy and triggered a conditional refinement. In the subsequent feedback-guided re-evaluation, the Rhetoric Agents revised their assessments, aligning with the Context Agent towards ”Not Sarcastic.” This revised consensus allowed the Decision Agent to correctly classify the input, demonstrating REAgent’s effectiveness in error correction and enhancing judgment robustness.
4.5.2 Quantitative Evaluation of Explanations
Furthermore, to quantitatively assess the utility of the intermediate reasoning generated by CAF-I’s specialized agents, an auxiliary experiment is conducted. This evaluation, detailed in Appendix D, confirmed that the agent-generated explanations provide valuable discriminative information, thereby enhancing the framework’s interpretability.
4.6 Inference Efficiency Analysis
To assess practical trade-offs, we analyzed CAF-I’s inference efficiency and detection performance against standard CoT and ToT baselines, all using the gpt-4o-mini backbone for fair comparison. This evaluation, detailed in Table 2, used a 400-sample subset from our four benchmarks.
Table 2 indicates CAF-I achieves the highest accuracy and Macro-F1 scores. Its average inference time of 9.67 seconds per sample is highly competitive, nearly identical to that of ToT at 9.64 seconds and only marginally more than standard CoT at 8.93 seconds, despite CAF-I’s sophisticated multi-agent architecture.
These results demonstrate CAF-I not only significantly outperforms both baselines in detection accuracy but does so without a substantial increase in computational overhead. This highlights the efficiency of CAF-I’s collaborative reasoning design, leveraging architectural depth for superior accuracy while maintaining practical inference speed.
Table 2: The average inference time (s/sample) and corresponding detection performance compared to baselines, using gpt-4o-mini on a 400-sample subset.
Method Acc. Ma-F1 Avg. Inf Time
CoT 63.50 62.48 8.93
ToT 68.75 66.44 9.64
CAF-I 72.50 71.03 9.67
5 CONCLUSION
This paper introduced CAF-I, an LLM-driven multi-agent framework for robust irony detection, integrating specialized context, semantic, and rhetoric agents with collaborative refinement. Experiments confirmed CAF-I’s state-of-the-art zero-shot performance over existing methods. Ablation studies validated its component necessity and architectural robustness, while also highlighting its interpretability and error correction.
Appendix A Detail Experiment Set Up
A.1 Datasets
We evaluate CAF-I’s efficacy and generalizability using four established sarcasm detection benchmarks:
  • IAC-V1 (Lukin & Walker, 2013): This dataset comprises comments from online political debates, offering examples of sarcasm embedded within argumentative discourse.
  • IAC-V2 (Oraby et al., 2017): An extension of IAC-V1, providing a larger, more diverse collection of sarcastic/non-sarcastic statements from similar online discussion forums.
  • MuSTARD (Castro et al., 2019): Sourced from popular television shows, MuSTARD provides conversational context for each utterance.
  • SemEval-2018 Task 3 (Van Hee et al., 2018): Consists of English tweets annotated for irony, representing challenges of social media text like brevity, informal language, and implicit contextual cues for ironic intent.
Table 3 summarizes detailed dataset statistics. Their diversity in domains and challenges enables robust framework evaluation.
Table 3: Overview of the benchmark datasets used for evaluating irony detection.
Dataset Year Size Avg. Length Domain Context
IAC-V1 2013 320 68 Debate No
IAC-V2 2016 1042 43 Debate No
MuSTARD 2019 784 14 Dialogue Yes
SemEval-2018 2018 183 14 Twitter No
A.2 Evaluation Metrics
We evaluate model performance using Accuracy for overall correctness and the Macro-F1 score as the primary metric, following standard practices in the field (Yao et al., 2024). Macro-F1 averages the F1-scores for ironic and non-ironic classes independently, providing a balanced measure suitable for potentially imbalanced datasets common in irony detection.
A.3 Experiment Details
The GPT-4o model serves as the LLM backbone for all agents within the proposed CAF-I framework, accessed via the official OpenAI API. Its selection ensures state-of-the-art baseline capabilities in natural language understanding and reasoning. To enhance reproducibility, a temperature setting of 0 was used for all API interactions.
A.4 Comparison Baselines
To contextualize CAF-I’s performance, we compare it against a comprehensive suite of baselines across three categories: LLM-based: Approaches from the SarcasmCue framework (Yao et al., 2024), including GPT-4o Zero-Shot and three advanced prompting strategies: Chain of Contradiction, Graph of Cues, and Bagging of Cues. Performance figures are cited from their reported results. Fine-tuned PLMs: Standard Pre-trained Language Models such as BERT-base-uncased (Devlin et al., 2019) and RoBERTa-base (Liu et al., 2019), fine-tuned on target datasets. Deep Learning Methods: Influential deep learning methods with task-specific architectures, namely MIARN (Tay et al., 2018), SAWS (Pan et al., 2020), and DC-Net (Liu et al., 2021). Results are sourced from literature with aligned evaluation settings (Qiu et al., 2024; Zhang et al., 2024; Tay et al., 2018; Xue et al., 2024; Pan et al., 2020).
Appendix B Case Study Illustration
This appendix section presents a detailed visual walkthrough of the case study discussed in Section 4.5.1. Figure 4 below illustrates the step-by-step multi-agent reasoning, collaborative reanalysis, and conditional refinement process undertaken by the CAF-I framework on an example input. This visualization aims to further clarify CAF-I’s interpretability and its mechanism for error correction, as highlighted in the main text, by showing how the agents interact and revise their judgments to reach a final, more accurate classification.
Figure 4: Illustration of the CAF-I framework’s step-by-step reasoning and refinement process on a input example.
Appendix C Detailed Robustness Evaluation with Different LLM Backbones
This section provides the detailed performance metrics for the CAF-I framework when implemented with alternative Large Language Model backbones, namely Qwen 2-7B and Llama 3-8B. The results, as summarized in Table 4, are compared against each LLM’s baseline Input-Output prompt performance. This data supports the discussion in Section 4.4 regarding CAF-I’s architectural robustness and its ability to enhance performance across different underlying models.
Table 4: Robustness evaluation of the CAF-I framework using different LLM backbones Qwen 2-7B, Llama 3-8B compared to the IO prompting. Baseline IO results are aligned with SarcasmCue (Yao et al., 2024). All scores are in %.
LLM Backbone IAC-V1 IAC-V2 SemEval-2018 MuSTARD
Acc. Ma-F1 Acc. Ma-F1 Acc. Ma-F1 Acc. Ma-F1
Qwen2-7b + IO 56.56 49.32 51.85 38.57 45.15 38.83 54.78 46.17
Qwen2-7b + CAF-I 71.85 71.19 67.08 66.39 81.63 81.25 72.60 73.80
Llama3-8b + IO 55.94 46.40 54.70 43.74 49.36 44.46 54.64 44.99
Llama3-8b + CAF-I 60.62 60.40 70.36 73.55 75.76 75.47 59.56 57.11
Appendix D Quantitative Evaluation of Agent Explanations
An auxiliary experiment on the SemEval-2018 dataset, using GPT-3.5 Turbo as a baseline, assessed the utility of intermediate reasoning from CAF-I’s analysis agents. We compared GPT-3.5 Turbo’s performance under two conditions: using only the original input text with a standard IO prompt, versus augmenting the input with a feature prompt concatenating stage 1 textual explanations from CAF-I’s Context, Semantic, and Rhetoric agents. Incorporating these agent explanations improved the baseline model’s Macro-F1 score from 68.4% to 70.2%. This performance gain indicates that CAF-I’s specialized agents generate valuable, discriminative reasoning, substantiating the framework’s effective internal processing and enhancing its interpretability.
Appendix E Future Work
Future research could focus on extending CAF-I’s application beyond irony detection to include other forms of emotional detection and broader nuanced linguistic phenomena, as well as adapting it for cross-lingual applications. Further work might also involve developing the framework for multimodal contexts, enhancing agent collaboration dynamics, and systematically evaluating its performance and efficiency with a wider array of Large Language Model backbones.